OpenStudy (cakemeister):

HELP! I'm making potential energy equivalent to kinetic energy, but I think my equation for finding "v". So far it's gh=1/2v^2.

OpenStudy (ookawaiioo):

PE = KE mgh = 1/2mv^2 (mass is canceled out) gh = 1/2 v^2 sqrt(2gh) = velocity

OpenStudy (cakemeister):

Thank you!!!

OpenStudy (ookawaiioo):

:D